Mark's Uncle Cleo, is sitting under a pear tree. A 0.35 kg pear falls from a heigt of 15.0 m above him. The pear collides with Cleo causing him to sink into the lawn chair he was sitting on. Cleo and the pear go down for 0.020 s until coming to a stop. If Cleo's mass is 50-kg (a) What speed does it hit his head with? , (b) What speed will he and the pear have after the collision?, (c) What impulse has been exerted by the lawn chair? (d) What force was exerted bythe lawn chair?


1
Expert's answer
2021-03-08T08:28:52-0500

a)


"v=\sqrt{2gh}=\sqrt{2(9.8)(15)}=17.1\frac{m}{s}"

b)


"mv=(m+M)u\\0.35(17.1)=(0.35+50)u\\u=0.12\frac{m}{s}"

c)


"I=0.35(17.1)=6.0\ Ns"

d)


"F=\frac{6}{0.02}=300\ N"